Mauro Products distributes a single product, a woven basket whose selling price is $22 per unit and whose variable expense is $16 per unit. The company's monthly fixed expense is $14,400 


Required:

 1. Calculate the company's break-even point in unit sales.

 2. Calculate the company's break-even point in dollar sales. (Do not round intermediate calculations.)

 3. If the company's fixed expenses increase by $600, what would become the new break-even point in unit sales? In dollar sales? (Do not round intermediate calculations.) 
Answer #1
Answer
1 Break even point (units) = Fixed cost / Contribution margin per unit
= 14400 / (22-16)
= 2400 Baskets
2 Break even sales in dollars = Fixed cost / Contribution margin ratio
= 14400 / (6/22)
= $ 52800
3a. Break even point (units) = Fixed cost / Contribution margin per unit
= 15000 / (22-16)
= 2500 Baskets
3b. Break even sales in dollars = Fixed cost / Contribution margin ratio
= 15000 / (6/22)
= $ 55000